Hubble Space Telescope Status Report
06.18.09
 
On June 15, 2009 at approximately 2:51 a.m. EDT, the HST Science Instrument Command and Data Handler (SI C&DH), installed during the recently completed servicing mission, began to send unexpected "zero" readings to the on-board HST 486 computer. Seconds later, the computer sent commands to put the SI C&DH and the HST science instruments into safe mode.

Because of the SI C&DH's anomalous condition no science instrument telemetry was being reported, and Power Distribution Unit (PDU) currents and voltages provided the only indication of science instrument status. These data indicated that the Science Instruments had not transitioned to their safe states – indicating that none of them had received the HST486's safing commands. Engineers quickly concluded that the SI C&DH was neither receiving or forwarding commands nor processing telemetry. Other than these components, all HST spacecraft systems were nominal throughout and after the event.

Beginning at 2:50 pm, attempts to send commands to the HST payload via the SI C&DH interface used for ground commanding were unsuccessful.

As planned, the next step taken was to power-cycle the SI C&DH – that is, to turn it off and then on again. The power cycle succeeded just before 4 pm, restoring fixed-format telemetry from all payload elements. This telemetry indicated that each of the science instruments was in the state to which it had been commanded prior to the SI C&DH anomaly.

Subsequently each science instrument was commanded into its safe mode by HST's flight controllers.

Around 8 pm Monday evening, Hubble engineers recovered the SI C&DH to 'normal mode.' The science instruments will remain in their safe configurations while investigation of Monday's anomaly continues. Hubble engineers are assessing the risk posture of the science instruments from potential future reoccurrences of this anomaly.

An Anomaly Review Board (ARB) will be convened to further evaluate the anomaly and make recommendations to the HST Program. We are presently looking at approximately a one-week delay to our planned Servicing Mission Observatory Verification (SMOV) activities.

The Observatory as a whole, including the new and repaired instruments, is in excellent shape and activities not requiring the use of the SI C&DH (Gyro and FGS calibrations for instance) are proceeding.